# PSS-BA: LiDAR Bundle Adjustment with Progressive Spatial Smoothing

### v1

@@ Abstract
Figure 1: LiDAR Bundle Adjustment with Progressive Spatial Smoothing. Taking noisy LiDAR frames as input, PSS-BA progressively smooths out the noisy data and updates LiDAR poses, which ensures both the convergence and accuracy of state estimation. Ultimately, it enables the simultaneous generation of accurate LiDAR frames and parameterized surfaces and facilitates 3D modeling.
footnotetext: This research is supported by the National Research Foundation, Singapore, under its Medium-Sized Center for Advanced Robotics Technology Innovation (CARTIN). All authors are with School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore 639798, 50 Nanyang Avenue. (E-mail: jianping.li@ntu.edu.sg, thienminh.nguyen@ntu.edu.sg, shyuan@ntu.edu.sg, elhxie@ntu.edu.sg)
I Introduction
Accurate 3D reconstruction of in GNSS-denied complex environment is a fundamental task in the fields of photogrammetry [1, 2, 3] and robotics [4, 5], in which the exported point clouds is the basis for many tasks such as facility inspection [6], building information modeling (BIM) [7], and robot navigation [8]. Although commercial mobile mapping systems (e.g., GeoSLAM, Leica Pegasus, etc.) equipped with expensive sensors have shown their powerful mapping ability over the past decade [9, 10], with recent developments in light-weight solid LiDAR technologies [11], research on improving the quality of point cloud collected by low-cost mapping systems is a hot topic in both academia and industry [12, 13].
In general, point cloud data obtained from a mobile mapping system are subjected to two sources of error: the pose error and the LiDAR measurement error. Most existing SLAM systems utilize some sequential registration scheme (e.g., ICP [14], NDT [15], features [16]) to merge the laser scans together, the incremental processing will cause the accumulated error in the pose estimates. The pose error can be corrected using some LiDAR bundle adjustment method [17] (which we will review in Section I-A). As for the LiDAR measurement error, it is mainly caused by the range measurement noise, which can be corrected using point cloud smoothing or map-centric optimization [18] in 3D space (a thorough review is given in Section I-B). Intuitively, we found that the LiDAR bundle adjustment and point cloud smoothing [19] processes can benefit each other. On one hand, even with the imperfect initial poses, the local shape prior of the noisy point clouds can still be recognized by applying smoothing kernels [19] in 3D space. The shape prior can be used as a more general factor in bundle adjustment than the planar features. On the other hand, the adjusted poses guarantee the global accuracy of the point clouds compared to only conducting point cloud smoothing in a local 3D space. The related works of LiDAR bundle adjustment and point cloud smoothing are reviewed as follows.
I-A LiDAR Bundle Adjustment
@@
I-B Point Cloud Smoothing
@@
Overall, to tackle the challenges of existing LiDAR bundle adjustment methods in complex environments, we propose PSS-BA. As shown in Fig. 1, the core idea of the proposed PSS-BA is that, even with the imperfect initial poses, we still can obtain a shape prior by applying multi-scale smooth kernels on the initial noisy point clouds. Then the multi-scale smooth kernels are used for pose correction progressively. The main contributions of the proposed method can be stated as:
@@
The rest of this paper is structured as follows. A detailed description of the proposed method is presented in Section II. The experiments are conducted on simulation and in-house datasets in Section III. Conclusion and future work are drawn in Section IV.
II Methodology
II-A Notations and System Overview
II-A1 Poses
@@
II-A3 Polynomial Smoothing Kernel
@@
II-A4 System Overview
The proposed PSS-BA includes two key modules, namely spatial smoothing (II-B) and poses adjustment (II-C), which are illustrated in Fig.2. Taking the noisy LiDAR frames as input and a preset large smoothing kernel width $\gamma$, the spatial smoothing module smooths the noisy point cloud at a coarse scale and obtains the polynomial coefficients. Then the pose adjustment module utilizes the extracted polynomial coefficients to correct the lidar frames. If the change of correction is smaller than a threshold, it could be regarded as convergence. Otherwise, the smoothing kernel width $\gamma$ is decreased to a finer scale $\gamma\leftarrow\gamma/k$ to conduct the spatial smoothing again. The iterative smoothing and adjustment strategy guarantees the convergence and accuracy of the LiDAR bundle adjustment.
@@ II-A4 System Overview
II-B Spatial Smoothing for LiDAR Factor Extraction
II-B1 Smoothing Kernel Sampling
@@
II-B2 Weighted Surface Fitting
@@
II-B3 Points Smoothing and Factor Association
@@
II-C Poses Adjustment using Polynomial Surface Constraints
II-C1 Jacobian calculation of polynomial residual
@@
$\sigma_{i,j}$ is correlated with $i^{th}$ and $j^{th}$ poses, namely $[\hat{R}_{i},\hat{t}_{i}]$ and $[\hat{R}_{j},\hat{t}_{j}]$ as shown in Fig. 3. The jacobian of $\sigma_{i,j}$ with respect to the $i^{th}$ pose can be derived using the chain rule as follows:
$\begin{split}&\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{t}}_{i}}=\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{j}^{S}}\frac{\partial\hat{\mathbf{p}}_{j}^{S}}{\partial\hat{\mathbf{p}}_{i}^{W}}\frac{\partial\hat{\mathbf{p}}_{i}^{W}}{\partial\hat{\mathbf{t}}_{i}}=-\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{i}^{S}}\hat{\mathbf{M}}_{i},\\ &\frac{\partial\sigma_{i,j}}{\partial\hat{\bm{\theta}}_{i}}=\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{j}^{S}}\frac{\partial\hat{\mathbf{p}}_{j}^{S}}{\partial\hat{\mathbf{p}}_{i}^{W}}\frac{\partial\hat{\mathbf{p}}_{i}^{W}}{\partial\hat{\bm{\theta}}_{i}}=\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{i}^{S}}\hat{\mathbf{M}}_{i}\left[\hat{\mathbf{R}}_{i}\mathbf{p}_{i}^{L}\right]_{\times},\\ &\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{j}^{S}}=\left[2\mathbf{\alpha}_{0}^{i}\hat{x}_{j}^{S}+\mathbf{\alpha}_{2}^{i}\hat{y}_{j}^{S}+\mathbf{\alpha}_{3}^{i},2\mathbf{\alpha}_{1}^{i}\hat{y}_{j}^{S}+\mathbf{\alpha}_{2}^{i}\hat{x}_{j}^{S}+\mathbf{\alpha}_{4}^{i},-1\right].\end{split}$ (11)
Similarly, the jacobian of $\sigma_{i,j}$ respect with the $j^{th}$ pose is as follows:
$\begin{split}&\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{t}}_{j}}=\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{j}^{S}}\frac{\partial\hat{\mathbf{p}}_{j}^{S}}{\partial\hat{\mathbf{p}}_{j}^{W}}\frac{\partial\hat{\mathbf{p}}_{j}^{W}}{\partial\hat{\mathbf{t}}_{j}}=\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{j}^{S}}\hat{\mathbf{M}}_{i},\\ &\frac{\partial\sigma_{i,j}}{\partial\hat{\bm{\theta}}_{j}}=\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{j}^{S}}\frac{\partial\hat{\mathbf{p}}_{j}^{S}}{\partial\hat{\mathbf{p}}_{j}^{W}}\frac{\partial\hat{\mathbf{p}}_{j}^{W}}{\partial\hat{\bm{\theta}}_{j}}=-\frac{\partial\sigma_{i,j}}{\partial\hat{\mathbf{p}}_{j}^{S}}\hat{\mathbf{M}}_{i}\left[\hat{\mathbf{R}}_{j}\mathbf{p}_{j}^{L}\right]_{\times}.\end{split}$ (12)
II-C2 Poses correction
@@ III-B Simulation Dataset
III-C Real-world Dataset
We collected a lot of in-house datasets using a helmet-based mapping system consisting of a Livox Mid360 laser scanner (Fig. 7) [1] in the Nanyang Technological University (NTU) campus. The initial values for the bundle adjustment are also obtained using Fast-LIO [4]. The laser scanning data are split into groups with 100 sequential frames. Some sample results from the proposed PSS-BA are shown in Fig. 8. From the visual inspection, the resulting point clouds from the proposed PSS-BA are very sharp and clear, even the indoor environments with irregular and complex 3D shapes, which may cause difficulty for the plane-based bundle adjustment methods.
Figure 7: An in-house helmet-based mapping system.
@@ III-C Real-world Dataset
IV Conclusion and future work
This paper proposes PSS-BA, targeting the accurate construction of point clouds from LiDAR data in complex environments. Existing solutions predominantly depend on detecting planar features, which may be inadequate in complex environments with less structural geometry or substantial initial pose errors. PSS-BA overcomes these challenges by combining spatial smoothing and poses adjustment modules, ensuring local consistency and global accuracy, leading to precise poses and high-quality point cloud reconstruction. PSS-BA’s effectiveness has been demonstrated in simulation and real-world datasets, offering a promising solution for accurate 3D modeling in challenging environments.
Due to the memory limitation, PSS-BA can only process several hundred frames together, which limits the PSS-BA’s application on large-scale mapping applications. In the near future, we will extend PSS-BA to large-scale applications by seeking hierarchical adjustment strategies.
Appendix for auxiliary optimization
To minimize Eq. (5), we use the auxiliary function $\bm{\Xi}$ introduced by [32]:
$\begin{split}\underset{\hat{\mathbf{n}},|\hat{\mathbf{n}}|=1,\bm{\Xi}}{\mathrm{argmin}}\ \mathbf{G}(\hat{\mathbf{n}}_{i})&=1-\hat{\mathbf{n}}^{\top}_{i}\breve{\mathbf{n}}_{i}+\beta|\mathbf{D}(\hat{\mathbf{n}}_{i})-\bm{\Xi}|^{2}+\mu|\bm{\Xi}|_{0},\end{split}$ (14)
where the $\beta$ controls how quikly the auxiliary function $\bm{\Xi}$ approaches Eq. (5), and is set to 0.01 at the begnining. With a initial guess of $\hat{\mathbf{n}}$, the above equation is optimized by fixing $\bm{\Xi}$ as follow:
$\begin{split}\underset{\bm{\Xi}}{\mathrm{argmin}}\ \mathbf{G}(\hat{\mathbf{n}}_{i})&=\beta|\mathbf{D}(\hat{\mathbf{n}}_{i})-\bm{\Xi}|^{2}+\mu|\bm{\Xi}|_{0},\end{split}$ (15)
The solution of Eq. (15) is given by:
[…]
Next, $\bm{\Xi}$ is fixed to optimize the normal $\hat{\mathbf{n}}$ as follow:
[…]
Eq. 17 is quadratic and can be optimized using least square estimation. It should be noted that as $|\hat{\mathbf{n}}|=1$, $\hat{\mathbf{n}}$ is updated using Eq. (3). The optimization of Eq. (15) and Eq. (17) iteratively with $\beta\leftarrow 2\beta$ to make $\bm{\Xi}$ approach $\mathbf{D}(\hat{\mathbf{n}}_{i})$.

### v2

@@ Abstract
footnotetext: This research was conducted under project WP5 within the Delta-NTU Corporate Lab with funding support from A*STAR under its IAF-ICP programme (Grant no: I2201E0013) and Delta Electronics Inc. This research is also supported by the National Research Foundation, Singapore, under its Medium-Sized Center for Advanced Robotics Technology Innovation (CARTIN). All authors are with School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore 639798, 50 Nanyang Avenue. (E-mail: jianping.li@ntu.edu.sg, thienminh.nguyen@ntu.edu.sg, shyuan@ntu.edu.sg, elhxie@ntu.edu.sg)
1 Introduction
Accurate 3D reconstruction of in GNSS-denied complex environment is a fundamental task in the fields of photogrammetry [1, 2, 3] and robotics [4, 5], in which the exported point clouds is the basis for many tasks such as facility inspection [6], building information modeling (BIM) [7], and robot navigation [8]. Although commercial mobile mapping systems (e.g., GeoSLAM, Leica Pegasus, etc.) equipped with expensive sensors have shown their powerful mapping ability over the past decade [9, 10], with recent developments in light-weight solid LiDAR technologies [11], research on improving the quality of point cloud collected by low-cost mapping systems is a hot topic in both academia and industry [12, 13, 14].
In general, point cloud data obtained from a mobile mapping system are subjected to two sources of error: the pose error and the LiDAR measurement error. Most existing SLAM systems utilize some sequential registration scheme (e.g., ICP [15], NDT [16], features [17]) to merge the laser scans together, the incremental processing will cause the accumulated error in the pose estimates. The pose error can be corrected using some LiDAR bundle adjustment method [18] (which we will review in Section 1.1). As for the LiDAR measurement error, it is mainly caused by the range measurement noise, which can be corrected using point cloud smoothing or map-centric optimization [19] in 3D space (a thorough review is given in Section 1.2). Intuitively, we found that the LiDAR bundle adjustment and point cloud smoothing [20] processes can benefit each other. On one hand, even with the imperfect initial poses, the local shape prior of the noisy point clouds can still be recognized by applying smoothing kernels [20] in 3D space. The shape prior can be used as a more general factor in bundle adjustment than the planar features. On the other hand, the adjusted poses guarantee the global accuracy of the point clouds compared to only conducting point cloud smoothing in a local 3D space. The related works of LiDAR bundle adjustment and point cloud smoothing are reviewed as follows.
1.1 LiDAR Bundle Adjustment
@@
1.2 Point Cloud Smoothing
@@
Overall, to tackle the challenges of existing LiDAR bundle adjustment methods in complex environments, we propose PSS-BA. The core idea of the proposed PSS-BA is that, even with the imperfect initial poses, we still can obtain a shape prior by applying multi-scale smooth kernels on the initial noisy point clouds. Then the multi-scale smooth kernels are used for pose correction progressively. The main contributions of the proposed method can be stated as:
@@
The rest of this paper is structured as follows. A detailed description of the proposed method is presented in Section 2. The experiments are conducted on simulation and in-house datasets in Section 3. Conclusion and future work are drawn in Section 4.
2 Methodology
2.1 Notations and System Overview
2.1.1 Poses
@@
2.1.3 Polynomial Smoothing Kernel
@@
2.1.4 System Overview
The proposed PSS-BA includes two key modules, namely spatial smoothing (2.2) and poses adjustment (2.3), which are illustrated in Fig.1. Taking the noisy LiDAR frames as input and a preset large smoothing kernel width $\gamma$, the spatial smoothing module smooths the noisy point cloud at a coarse scale and obtains the polynomial coefficients. Then the pose adjustment module utilizes the extracted polynomial coefficients to correct the lidar frames. If the change of correction is smaller than a threshold, it could be regarded as convergence. Otherwise, the smoothing kernel width $\gamma$ is decreased to a finer scale $\gamma\leftarrow\gamma/k$ to conduct the spatial smoothing again. The iterative smoothing and adjustment strategy guarantees the convergence and accuracy of the LiDAR bundle adjustment.
@@ 2.1.4 System Overview
2.2 Spatial Smoothing for LiDAR Factor Extraction
2.2.1 Smoothing Kernel Sampling
@@
2.2.2 Weighted Surface Fitting
@@
2.2.3 Points Smoothing and Factor Association
@@
2.3 Poses Adjustment using Polynomial Surface Constraints
2.3.1 Jacobian calculation of polynomial residual
@@
$\sigma_{i,j}$ is correlated with $i^{th}$ and $j^{th}$ poses, namely $[\hat{R}_{i},\hat{t}_{i}]$ and $[\hat{R}_{j},\hat{t}_{j}]$ as shown in Fig. 2. The jacobian of $\sigma_{i,j}$ with respect to the $i^{th}$ pose can be derived using the chain rule.
2.3.2 Poses correction
@@ 3.2 Simulation Dataset
3.3 Real-world Dataset
We collected a lot of in-house datasets using a helmet-based mapping system consisting of a Livox Mid360 laser scanner [2] in the Nanyang Technological University (NTU) campus. The initial values for the bundle adjustment are also obtained using Fast-LIO [4]. The laser scanning data are split into groups with 100 sequential frames. Some sample results from the proposed PSS-BA are shown in Fig. 6. From the visual inspection, the resulting point clouds from the proposed PSS-BA are very sharp and clear, even the indoor environments with irregular and complex 3D shapes, which may cause difficulty for the plane-based bundle adjustment methods.
@@ 3.3 Real-world Dataset
4 Conclusion
This paper proposes PSS-BA, targeting the accurate construction of point clouds from LiDAR data in complex environments. Existing solutions predominantly depend on detecting planar features, which may be inadequate in complex environments with less structural geometry or substantial initial pose errors. PSS-BA overcomes these challenges by combining spatial smoothing and pose adjustment modules, ensuring local consistency and global accuracy, leading to precise poses and high-quality point cloud reconstruction. PSS-BA’s effectiveness has been demonstrated in simulation and real-world datasets, offering a promising solution for accurate 3D modeling in challenging environments.